4 0 Required information Problem 05.028 DEPENDENT MULTI-PART PROBLEM . ASSIGN ALL PARTS Parker Hennifin of Cleveland, Ohio manufactures CNG fuel dispensers It needs replacement equipment to streamline one of its production lines for a new contract, but it plans to sell the equipment at or before iEs expected ife is reached at an estimabed market value for used equlpment Problem 05.028.a Future Worth Analysis Select between the two options using te corporate MARR of 15% per year and a fare wort, analysis for the expected use period. The future worth of option D is S The fufure worth of option E i Option (ak o seec is selected /2
Answer #1

FV of D = -FV(15%, 6, -20000, -80000) - FV(15%, 3, (-80000+11500)) + 11500 = -586,485.88

FV of E = =-FV(15%, 6, -16000, -90000) + 12500 = -335,735.28

E is better option.